OpenStudy (anonymous):

how many quarts of water do you need to add to 40 quarts of 80 percent alcohol to get a solution of 10 percent alcohol?

OpenStudy (anonymous):

\[ 40\cdot 0.2+x=(x+40)\cdot 0.9 \] Solve that.

OpenStudy (sriram):

vol of alcohol=40*80/100=32 vol of final soln such that alc=10% means 320quarts hence water to be added=320-40=280

OpenStudy (anonymous):

Those two answers are equivalent.

OpenStudy (anonymous):

fyi